how many grams of NaoH are required to fully consume 3.10 grams of H2so4


1
Expert's answer
2021-03-02T04:56:26-0500
"n(H2SO4)=3.198=0.0315 mol"

2NaOH+H2SO4=2H2O+Na2SO4


"n(NaOH)\/2=0.0315\/1"

"n(NaOH)=0.0315*2=0.063"

"m(NaOH)=nM=0.063*40=2.52 g"